### v1

@@ 3.2 Pareto-optimal Trade-offs
The trade-off between communication and computation is illustrated in Fig. 2, where each gray point reflects a specific choice of $\alpha$ and $\beta$ in the range from 1 to 100. The coordinates of each point represent the estimated number of communication steps (horizontal axis) and number of computation steps (vertical axis) required to achieve an accuracy of $\epsilon$, obtained from Table 2. It can be observed that, according to the definition of a Pareto-optimal solution in multi-objective optimization, a specific set of parameters lies on the Pareto frontier, as shown by the bold red lines. This set achieves the lowest overall complexity for reaching a given accuracy level, meaning that improvements in either communication or computation complexity are not possible without compromising the other.
Definition 1 (Pareto-optimality [4, 7]).
For a multi-objective optimization problem that minimizes $\left\{g_{i}\left(\omega\right)\right\}_{j\in\left[m\right]}$, a point $\omega$ is Pareto-optimal if there does not exist another point $\omega^{\prime}$ such that: $\forall j\in[m]$, $g_{j}\left(\omega^{\prime}\right)\leqslant g_{j}\left(\omega\right)$; and $\exists j\in\left[m\right]$, $g_{j}\left(\omega^{\prime}\right)<g_{j}\left(\omega\right)$. The set of Pareto-optimal points is called the Pareto frontier.
The following corollary shows that by appropriately incorporating accelerated gossip communication, Acc-FlexGT substantially improves the iteration complexity and achieves Pareto-optimal trade-offs between communication and computation.
Corollary 2.
Suppose Assumptions 24 hold. Let the stepsize satisfy the condition (9) given in Theorem 1, and denote by $\hat{K}$ the number of iterations required for Acc-FlexGT to reach an arbitrary accuracy $\epsilon>0$.
@@
(16) $\alpha=\lceil\frac{\max\left\{\ln 2,1/2\ln\left(n\beta L/\mu\right)\right\}}{\sqrt{1-\sqrt{\rho_{W}}}}\rceil,$
@@
(17) $\hat{K}=\mathcal{O}\left(\frac{L}{\mu}\log\frac{\left[V_{0}\right]}{\epsilon}+\frac{\sigma^{2}}{n\mu^{2}\beta\epsilon}\right).$
@@
(18) $\alpha=\lceil\frac{\max\left\{\ln 2,1/2\ln\left(n\beta\right)\right\}}{\sqrt{1-\sqrt{\rho_{W}}}}\rceil,$
@@
(19) $\hat{K}=\mathcal{O}\left(\frac{\sqrt{L}}{\epsilon}+\frac{L\sigma^{2}}{n\beta\epsilon^{2}}\right)\mathbb{E}\left[V_{0}\right].$
@@
(20) $\alpha=\lceil{\frac{\max\left\{\ln 2,1/2\ln\left(\beta\max\left\{n,R_{0}\right\}\right)\right\}}{\sqrt{1-\sqrt{\rho_{W}}}}}\rceil$
@@
(21) $\hat{K}=\mathcal{O}\left(\frac{L}{\epsilon}+\frac{L\sigma^{2}}{n\beta\epsilon^{2}}\right)\mathbb{E}\left[f\left(\bar{x}_{0}\right)-f^{*}\right].$
Proof.
The proof is similar to that of Corollary 1. By Lemma 1, we have $1-\bar{\rho}_{W}\geqslant 1/2$, and the topology effect part $\varUpsilon_{2}$ can be dominated when $\bar{\rho}_{W}$ is sufficiently small. Specifically, let $\hat{\gamma}$ denote the obtained upper bound of the stepsize in (9), and let the stepsize $\gamma$ satisfy
$\min\left\{\frac{\ln\left(\max\left\{2,\beta^{2}\mu^{2}\hat{K}/\varUpsilon_{1}\right\}\right)}{\beta\mu\hat{K}},\,\,\hat{\gamma}\right\}$
for the strongly convex case; and
$\min\left\{\frac{1}{\sqrt{\varUpsilon_{1}\hat{K}}},\,\,\hat{\gamma}\right\}$
Based on Corollary 2, we derive the communication and computation complexities of Acc-FlexGT, as shown in Table 3. Compared to FlexGT, Acc-FlexGT achieves a Pareto-optimal trade-off in terms of $\beta$ by selecting specific values of $\alpha$, as outlined in Corollary 2. In other words, there is no choice of $\beta$ that can simultaneously reduce both communication and computation complexity. This distinguishes our work from existing methods with accelerated communication [31, 10], which generally do not consider both communication and computational efficiency simultaneously, nor the trade-offs between them.
Remark 3.
Compared to existing results, Acc-FlexGT recovers DeTAG [31] as a special case when $\alpha=\beta$. Moreover, the iteration complexity (21) obtained for the nonconvex case is optimal in the sense that it matches the lower bound established in [31, 54], up to a logarithmic factor arising from the choice of $\beta=\alpha=\tilde{\mathcal{O}}\left(1/\sqrt{1-\sqrt{\rho_{W}}}\right)$. In the strongly convex case, our result improves the first term in (17) from $1/\sqrt{\epsilon}$ to $\log\left(1/\epsilon\right)$ compared to [10], as we use a constant stepsize instead of a stepsize that decays with iterations. To the best of our knowledge, the convex case is first considered in this work.
@@ 4 Convergence Analysis
Proof sketch. The basic idea of the proof is illustrated in Fig. 3, which builds on the established contraction properties of the optimality gap for (strongly) convex objective functions (Lemma 2) and descent property for the nonconvex case (Lemma 3), both of which involve dependencies on the consensus and GT errors. Rather than using a direct recursive analysis, the key to bound these errors is to break the cyclic dependency between the consensus and GT errors (Lemmas 5 and 6) as highlighted by the arrows in the figure. For the (strongly) convex case, we achieve this by designing a suitable Lyapunov function in (8). Conversely, for the nonconvex case, we accomplish this through a straightforward method of accumulating the errors (Lemmas 7 and 8). Notably, the proof relies on a double-loop analysis. We first derive the upper bounds of the client divergence over the interval between two communications (inner loop) (cf., Lemma 4 in Appendix A). These bounds are subsequently utilized to establish the contraction properties of the consensus and GT errors during each communication round (outer loop). Following this proof sketch, we begin by introducing some key lemmas that are crucial for the proof of Theorem 1.
@@ 4.1 Key Lemmas
For the (strongly) convex case, i.e., Assumption 1 hold, we have the following contraction property for the optimality gap.
Lemma 2 (Optimality gap).
Suppose Assumptions 14 hold. Let the step-size satisfy $\gamma\leqslant\frac{1}{4\beta L}$. Then, we have for all $k\geqslant 0$,
(22) $\mathbb{E}\left[\left\|\bar{x}_{\beta\left(k+1\right)}-x^{*}\right\|^{2}\right]$ $\leqslant\left(1-\frac{\mu\beta\gamma}{2}\right)\mathbb{E}\left[\left\|\bar{x}_{\beta k}-x^{*}\right\|^{2}\right]+\frac{\gamma^{2}\beta\sigma^{2}}{n}$
@@
(23) $\mathbb{E}\left[\left\|\bar{x}_{\beta\left(k+1\right)}-x^{*}\right\|^{2}\right]$ $=\mathbb{E}\left[\left\|\bar{x}_{\beta k}-x^{*}\right\|^{2}\right]+\gamma^{2}\underset{=:S_{2}}{\underbrace{\mathbb{E}\left[\left\|\sum_{j=0}^{\beta-1}{\frac{\mathbf{1}^{\top}}{n}\mathbf{y}_{\beta k+j}}\right\|^{2}\right]}}$
$\quad-2\gamma\beta\underset{=:S_{3}}{\underbrace{\mathbb{E}\left[\left<\bar{x}_{\beta k}-x^{*},\frac{\mathbf{1}^{\top}}{n}\nabla F_{\beta k}\right>\right]}},$
where we used the fact that $\mathbb{E}\left[\nabla G_{\beta k}\right]=\nabla F_{\beta k}$.
@@
(24) $\mathbb{E}\left[\left\|\sum_{j=0}^{\beta-1}{\frac{\mathbf{1}^{\top}}{n}\mathbf{y}_{\beta k+j}}\right\|^{2}\right]$ $\leqslant\frac{\beta\sigma^{2}}{n}+\mathbb{E}\left[\left\|\sum_{j=0}^{\beta-1}{\frac{1}{n}\sum_{i=1}^{n}{\nabla f_{i}\left(x_{i,\beta k}\right)}}\right\|^{2}\right]$
@@
where we have used Assumption 3 on bounded stochastic gradient variance and Assumption 2 on smoothness of the objectives. For $S_{3}$, using the convexity and smoothness of $f_{i}$ in Assumption 1 and 2, we have
@@
$=\frac{1}{n}\sum_{i=1}^{n}{\mathbb{E}\left[\left<x_{i,\beta k}-x^{*},\nabla f_{i}\left(x_{i,\beta k}\right)\right>\right]}-\frac{1}{n}\sum_{i=1}^{n}{\mathbb{E}\left[\left<x_{i,\beta k}-\bar{x}_{\beta k},\nabla f_{i}\left(x_{i,\beta k}\right)\right>\right]}$
$\geqslant\mathbb{E}\left[f\left(\bar{x}_{\beta k}\right)-f\left(x^{*}\right)\right]+\frac{\mu}{2n}\mathbb{E}\left[\left\|\mathbf{x}_{\beta k}-\mathbf{1}x^{*}\right\|^{2}\right]-\frac{L}{2n}\mathbb{E}\left[\left\|\tilde{\mathbf{x}}_{\beta k}\right\|^{2}\right].$
@@
(26) $-\mathbb{E}\left[\left\|\mathbf{x}_{\beta k}-\mathbf{1}x^{*}\right\|^{2}\right]$
$=-\mathbb{E}\left[\left\|\tilde{\mathbf{x}}_{\beta k}\right\|^{2}\right]-n\mathbb{E}\left[\left\|\bar{x}_{\beta k}-x^{*}\right\|^{2}\right]-2\sum_{i=1}^{n}{\left(\mathbb{E}\left[\left<x_{i,\beta k}-\bar{x}_{\beta k},\bar{x}_{\beta k}-x^{*}\right>\right]\right)}$
@@
(27) $\mathbb{E}\left[\left\|\bar{x}_{\beta\left(k+1\right)}-x^{*}\right\|^{2}\right]$
@@
$\quad+\left(\frac{2\gamma^{2}\beta^{2}L^{2}}{n}+\frac{2\gamma\beta L}{n}\right)\mathbb{E}\left[\left\|\tilde{\mathbf{x}}_{\beta k}\right\|^{2}\right]-\left(2\gamma\beta-4\gamma^{2}\beta^{2}L\right)\mathbb{E}\left[f\left(\bar{x}_{\beta k}\right)-f\left(x^{*}\right)\right].$
Then, letting the stepsize satisfy $\gamma\leqslant\frac{1}{4\beta L}$, we complete the proof.
Next, we introduce lemmas for the nonconvex case, which play a key role in proving the convergence result in (12).
Lemma 3 (Descent lemma).
@@
(28) $\frac{1}{K}\sum_{k=0}^{K-1}{\mathbb{E}\left[\left\|\nabla f\left(\bar{x}_{\beta k}\right)\right\|^{2}\right]}$
$\leqslant\frac{4\left(\mathbb{E}\left[f\left(\bar{x}_{0}\right)\right]-\mathbb{E}\left[f^{*}\right]\right)}{\gamma\beta K}+\frac{4L^{2}}{n}\frac{1}{K}\sum_{k=0}^{K-1}{\mathbb{E}\left[\left\|\tilde{\mathbf{x}}_{\beta k}\right\|^{2}\right]}+\frac{2\gamma L}{n}\sigma^{2}.$
Proof.
@@
(29) $f\left(\bar{x}_{\beta\left(k+1\right)}\right)$ $\leqslant f\left(\bar{x}_{\beta k}\right)+\left<\nabla f\left(\bar{x}_{\beta k}\right),\bar{x}_{\beta\left(k+1\right)}-\bar{x}_{\beta k}\right>+\frac{L}{2}\left\|\bar{x}_{\beta\left(k+1\right)}-\bar{x}_{\beta k}\right\|^{2}.$
Then, for the inner-product term on the right, by the recursion of decision variable $x$ in (3a) and noticing that $\mathbb{E}\left[\nabla G_{\beta k+j}\right]=\nabla F_{\beta k}$, we have
@@
$\leqslant-\frac{\gamma}{2}\beta\mathbb{E}\left[\left\|\nabla f\left(\bar{x}_{\beta k}\right)\right\|^{2}\right]+\frac{\gamma}{2}\mathbb{E}\left[\left\|\sum_{j=0}^{\beta-1}{\frac{\mathbf{1}^{\top}}{n}\left(\nabla F_{\beta k}-\nabla F\left(\mathbf{1}\bar{x}_{\beta k}\right)\right)}\right\|^{2}\right]$
@@
(31) $\mathbb{E}\left[f\left(\bar{x}_{\beta\left(k+1\right)}\right)\right]$ $\leqslant\mathbb{E}\left[f\left(\bar{x}_{\beta k}\right)\right]-\left(\frac{\gamma\beta}{2}-\gamma^{2}\beta^{2}L\right)\mathbb{E}\left[\left\|\nabla f\left(\bar{x}_{\beta k}\right)\right\|^{2}\right]$
@@
(32) $\mathbb{E}\left[\left\|\nabla f\left(\bar{x}_{\beta k}\right)\right\|^{2}\right]\leqslant\frac{4\mathbb{E}\left[f\left(\bar{x}_{\beta k}\right)-f\left(\bar{x}_{\beta\left(k+1\right)}\right)\right]}{\gamma\beta}+\frac{4L^{2}}{n}\mathbb{E}\left[\left\|\tilde{\mathbf{x}}_{\beta k}\right\|^{2}\right]+\frac{2\gamma L}{n}\sigma^{2}.$

### v2

@@ 3.2 Pareto-optimal Trade-offs
The trade-off between communication and computation is illustrated in Fig. 2, where each gray point reflects a specific choice of $\alpha$ and $\beta$ in the range from 1 to 100. The coordinates of each point represent the estimated number of communication steps (horizontal axis) and number of computation steps (vertical axis) required to achieve an accuracy of $\epsilon$, obtained from Table 2. It can be observed that, under Definition 1 of Pareto-optimality in multi-objective optimization, a specific set of parameters lies on the Pareto frontier, as shown by the bold red lines. This set achieves the lowest overall complexity for reaching a given accuracy level, meaning that improvements in either communication or computational complexity are not possible without compromising the other.
The following corollary shows that by incorporating accelerated gossip communication, Acc-FlexGT substantially improves the iteration complexity and achieves Pareto-optimal trade-offs between communication and computation.
Corollary 2
Suppose Assumptions 24 hold. Let the stepsize satisfy the condition (9) given in Theorem 1, and denote by $\hat{K}$ the number of rounds required for Acc-FlexGT to reach an arbitrary accuracy $\epsilon>0$. Then,
@@
$\alpha=\lceil\frac{\max\left\{\ln 2,1/2\ln\left(n\beta L/\mu\right)\right\}}{\sqrt{1-\sqrt{\rho_{W}}}}\rceil,$ (16)
@@
$\hat{K}=\mathcal{O}\left(\frac{L}{\mu}\log\frac{\left[V_{0}\right]}{\epsilon}+\frac{\sigma^{2}}{n\mu^{2}\beta\epsilon}\right).$ (17)
@@
$\alpha=\lceil\frac{\max\left\{\ln 2,1/2\ln\left(n\beta\right)\right\}}{\sqrt{1-\sqrt{\rho_{W}}}}\rceil,$ (18)
@@
$\hat{K}=\mathcal{O}\left(\frac{\sqrt{L}}{\epsilon}+\frac{L\sigma^{2}}{n\beta\epsilon^{2}}\right)\mathbb{E}\left[V_{0}\right].$ (19)
@@
$\alpha=\lceil{\frac{\max\left\{\ln 2,1/2\ln\left(\beta\max\left\{n,R_{0}\right\}\right)\right\}}{\sqrt{1-\sqrt{\rho_{W}}}}}\rceil$ (20)
@@
$\hat{K}=\mathcal{O}\left(\frac{L}{\epsilon}+\frac{L\sigma^{2}}{n\beta\epsilon^{2}}\right)\mathbb{E}\left[f\left(\bar{x}_{0}\right)-f^{*}\right].$ (21)
Proof 3.3.
See Section 4.4
Based on Corollary 2, the communication and computational complexities of Acc-FlexGT are obtained by multiplying $\hat{K}$ by $\alpha$ and $\beta$, respectively, as summarized in Table 3. Compared to FlexGT, Acc-FlexGT achieves Pareto-optimal trade-offs with respect to $\beta$ by properly selecting $\alpha$, in the sense that it attains the Pareto frontier of the multi-objective optimization problem that jointly minimizes communication and computational costs. In other words, any reduction in communication complexity entails an increase in computational complexity, and vice versa, as indicated by the red curves in Fig. 2. This distinguishes our work from existing methods with accelerated communication [30, 54, 10], which achieve optimal iteration complexity but do not explicitly account for the communication and computational costs incurred per iteration. Instead, we provide a more flexible framework to achieve Pareto-optimal trade-offs between these two costs via the tunable parameters $\alpha$ and $\beta$, and recover the results of the aforementioned methods as special cases, as detailed below.
Remark 3
Compared to existing results, for the nonconvex case, the obtained computational complexity of Acc-FlexGT
$\tilde{\mathcal{O}}\left(\frac{L\sigma^{2}}{n\epsilon^{2}}+\frac{L}{\sqrt{1-\sqrt{\rho_{W}}}\epsilon}\right)$ (22)
is optimal in the sense that it matches the lower bound on the iteration complexity in [30, 54], up to logarithmic factors arising from the choice of
$\beta=\alpha=\tilde{\mathcal{O}}\left(\frac{1}{\sqrt{1-\sqrt{\rho_{W}}}}\right).$ (23)
Moreover, Acc-FlexGT achieves the optimal communication complexity
$\tilde{\mathcal{O}}\left(\frac{L}{\sqrt{1-\sqrt{\rho_{W}}}\epsilon}\right),$ (24)
with a sufficient number of local updates, i.e.,
$\beta\geqslant\lceil\frac{\sigma^{2}}{n\epsilon}\rceil,$ (25)
matching the lower bounds established for the deterministic case in [44, 54]. In the strongly convex case, our result improves the first term in (17) from $1/\sqrt{\epsilon}$ to $\log\left(1/\epsilon\right)$ compared to [10], as we use a constant stepsize instead of a stepsize that decays with iterations. To the best of our knowledge, in the convex case, this work is the first to establish results for GT methods with local updates and Chebyshev acceleration.
@@ 4 Convergence Analysis
Proof sketch. The basic idea of the proof is illustrated in Fig. 3, which builds on the established contraction properties of the optimality gap for (strongly) convex objective functions (Lemma 2) and descent property for the nonconvex case (Lemma 3), both of which involve dependencies on the consensus and GT errors. Rather than using a direct recursive analysis, the key to bounding these errors is to break the cyclic dependency between the consensus and GT errors (Lemmas 5 and 6) as highlighted by the arrows in Fig. 3. For the (strongly) convex case, we achieve this by designing a suitable Lyapunov function in (8). Conversely, for the nonconvex case, we accomplish this through a straightforward method of accumulating the errors (Lemmas 7 and 8). Notably, the proof relies on a double-loop analysis. We first derive the upper bounds of the client divergence over the interval between two communications (inner loop) (cf., Lemma 4 in the appendix). These bounds are subsequently utilized to establish the contraction properties of the consensus and GT errors during each communication round (outer loop). Following this proof sketch, we begin by introducing some key lemmas that are crucial for the proof of Theorem 1.
@@ 4.1 Key Lemmas
For the (strongly) convex case, i.e., when Assumption 1 holds, we obtain the following lemma for the optimality gap.
Lemma 2
$\mathbb{E}\left[\left\|\bar{x}_{\beta\left(k+1\right)}-x^{*}\right\|^{2}\right]$ (26)
@@
$\bar{x}_{\beta\left(k+1\right)}=\bar{x}_{\beta k}-\gamma\sum_{j=0}^{\beta-1}{\frac{\mathbf{1}^{\top}}{n}\mathbf{y}_{\beta k+j}}.$
Noticing that $\mathbf{1}^{\top}\bar{W}=\mathbf{1}^{\top}$ and the fact that $\mathbb{E}\left[G_{\beta k}\right]=\nabla F_{\beta k}$, we get
$\mathbb{E}\left[\left\|\bar{x}_{\beta\left(k+1\right)}-x^{*}\right\|^{2}\right]$ (27)
$=\mathbb{E}\left[\left\|\bar{x}_{\beta k}-\gamma\sum_{j=0}^{\beta-1}{\frac{\mathbf{1}^{\top}}{n}\mathbf{y}_{\beta k+j}}-x^{*}\right\|^{2}\right]$
$=\mathbb{E}\left[\left\|\bar{x}_{\beta k}-x^{*}\right\|^{2}\right]+\gamma^{2}\underset{=:S_{2}}{\underbrace{\mathbb{E}\left[\left\|\sum_{j=0}^{\beta-1}{\frac{\mathbf{1}^{\top}}{n}\mathbf{y}_{\beta k+j}}\right\|^{2}\right]}}$
$\quad-2\gamma\beta\underset{=:S_{3}}{\underbrace{\mathbb{E}\left[\left<\bar{x}_{\beta k}-x^{*},\frac{\mathbf{1}^{\top}}{n}\nabla F_{\beta k}\right>\right]}}.$
@@
$\mathbb{E}\left[\left\|\sum_{j=0}^{\beta-1}{\frac{\mathbf{1}^{\top}}{n}\mathbf{y}_{\beta k+j}}\right\|^{2}\right]$ (28)
$\leqslant\frac{\beta\sigma^{2}}{n}+\mathbb{E}\left[\left\|\sum_{j=0}^{\beta-1}{\frac{1}{n}\sum_{i=1}^{n}{\nabla f_{i}\left(x_{i,\beta k}\right)}}\right\|^{2}\right]$
@@
where we have used Assumption 3 on bounded stochastic gradient variance and Assumption 2 on smoothness of the objectives. For $S_{3}$, using the convexity and smoothness of $f_{i}$ in Assumptions 1 and 2, we have
@@
$=\frac{1}{n}\sum_{i=1}^{n}{\mathbb{E}\left[\left<x_{i,\beta k}-x^{*},\nabla f_{i}\left(x_{i,\beta k}\right)\right>\right]}$
$\quad-\frac{1}{n}\sum_{i=1}^{n}{\mathbb{E}\left[\left<x_{i,\beta k}-\bar{x}_{\beta k},\nabla f_{i}\left(x_{i,\beta k}\right)\right>\right]}$
$\geqslant\mathbb{E}\left[f\left(\bar{x}_{\beta k}\right)-f\left(x^{*}\right)\right]+\frac{\mu}{2n}\mathbb{E}\left[\left\|\mathbf{x}_{\beta k}-\mathbf{1}x^{*}\right\|^{2}\right]$
$\quad-\frac{L}{2n}\mathbb{E}\left[\left\|\tilde{\mathbf{x}}_{\beta k}\right\|^{2}\right].$
@@
$-\mathbb{E}\left[\left\|\mathbf{x}_{\beta k}-\mathbf{1}x^{*}\right\|^{2}\right]$ (30)
$=-\mathbb{E}\left[\left\|\tilde{\mathbf{x}}_{\beta k}\right\|^{2}\right]-n\mathbb{E}\left[\left\|\bar{x}_{\beta k}-x^{*}\right\|^{2}\right]$
$\quad-2\sum_{i=1}^{n}{\left(\mathbb{E}\left[\left<x_{i,\beta k}-\bar{x}_{\beta k},\bar{x}_{\beta k}-x^{*}\right>\right]\right)}$
@@
$\mathbb{E}\left[\left\|\bar{x}_{\beta\left(k+1\right)}-x^{*}\right\|^{2}\right]$ (31)
@@
$\quad+\left(\frac{2\gamma^{2}\beta^{2}L^{2}}{n}+\frac{2\gamma\beta L}{n}\right)\mathbb{E}\left[\left\|\tilde{\mathbf{x}}_{\beta k}\right\|^{2}\right]$
$\quad-\left(2\gamma\beta-4\gamma^{2}\beta^{2}L\right)\mathbb{E}\left[f\left(\bar{x}_{\beta k}\right)-f\left(x^{*}\right)\right].$
Then, letting the stepsize $\gamma\leqslant\frac{1}{4\beta L}$, we complete the proof.
Next, we present the descent lemma for the gradient norm in the nonconvex case, which plays a key role in establishing the convergence result in (12).
Lemma 3
@@
$\frac{1}{K}\sum_{k=0}^{K-1}{\mathbb{E}\left[\left\|\nabla f\left(\bar{x}_{\beta k}\right)\right\|^{2}\right]}$ (32)
$\leqslant\frac{4\left(\mathbb{E}\left[f\left(\bar{x}_{0}\right)\right]-\mathbb{E}\left[f^{*}\right]\right)}{\gamma\beta K}$
$\quad+\frac{4L^{2}}{n}\frac{1}{K}\sum_{k=0}^{K-1}{\mathbb{E}\left[\left\|\tilde{\mathbf{x}}_{\beta k}\right\|^{2}\right]}+\frac{2\gamma L}{n}\sigma^{2}.$
Proof 4.2.
@@
$f\left(\bar{x}_{\beta\left(k+1\right)}\right)$ $\leqslant f\left(\bar{x}_{\beta k}\right)+\left<\nabla f\left(\bar{x}_{\beta k}\right),\bar{x}_{\beta\left(k+1\right)}-\bar{x}_{\beta k}\right$> (33)
$\quad+\frac{L}{2}\left\|\bar{x}_{\beta\left(k+1\right)}-\bar{x}_{\beta k}\right\|^{2}.$
Then, for the inner-product term on the right, by the recursion of decision variable $x$ in (3a) and noticing that $\mathbb{E}\left[G_{\beta k+j}\right]=\nabla F_{\beta k}$, we have
@@
$\leqslant-\frac{\gamma}{2}\beta\mathbb{E}\left[\left\|\nabla f\left(\bar{x}_{\beta k}\right)\right\|^{2}\right]$
$\quad+\frac{\gamma}{2}\mathbb{E}\left[\left\|\sum_{j=0}^{\beta-1}{\frac{\mathbf{1}^{\top}}{n}\left(\nabla F_{\beta k}-\nabla F\left(\mathbf{1}\bar{x}_{\beta k}\right)\right)}\right\|^{2}\right]$
@@
$\mathbb{E}\left[f\left(\bar{x}_{\beta\left(k+1\right)}\right)\right]$ (35)
$\leqslant\mathbb{E}\left[f\left(\bar{x}_{\beta k}\right)\right]-\left(\frac{\gamma\beta}{2}-\gamma^{2}\beta^{2}L\right)\mathbb{E}\left[\left\|\nabla f\left(\bar{x}_{\beta k}\right)\right\|^{2}\right]$
@@
$\mathbb{E}\left[\left\|\nabla f\left(\bar{x}_{\beta k}\right)\right\|^{2}\right]$ $\leqslant\frac{4\mathbb{E}\left[f\left(\bar{x}_{\beta k}\right)-f\left(\bar{x}_{\beta\left(k+1\right)}\right)\right]}{\gamma\beta}$ (36)
$\quad+\frac{4L^{2}}{n}\mathbb{E}\left[\left\|\tilde{\mathbf{x}}_{\beta k}\right\|^{2}\right]+\frac{2\gamma L}{n}\sigma^{2}.$